OpenStudy (anonymous):

Which of the following is the equation of the circle with center (4, 0) and r =√5. Choose one answer. a. x^2 + (y - 4)^2 = 5 b. x^2 + (y + 4)^2 =√5 c. (x + 4)^2 + y2 =√5 d. (x-4)^2+y^2=5

OpenStudy (amistre64):

\[(x-C_x)^2+(y-C_y)^2=r^2\]

OpenStudy (amistre64):

where Cx and Cy are the center parts

OpenStudy (anonymous):

(x-a)^2+(y-b)^2=r^2 where (a,b) is the centre of the circle and r is the radius

OpenStudy (anonymous):

are you not getting??